There are two unknown gas. Let's use gas X and Y. X is 0.68 times as fast as Y. The molar mass of Y is 17.0 g/mol. What is the molar mass of X?


1
Expert's answer
2021-02-11T04:49:02-0500

V1/V2 = √(m2/m1)

X/Y = 0.68

m1=m2/(X/Y)^2 = 17/0.68^2= 36,7647059.